Bergan Company estimates that total factory overhead costs will be $620,000 for the year. Direct labor hours are estimated to be 80,000.

Required:

a. For Bergan Company, determine the predetermined factory overhead rate using direct labor hours as the activity base.
b. During May, Bergan Company accumulated 2,500 hours of direct labor costs on Job 200 and 3,000 hours on Job 305. Determine the amount of factory overhead applied to Jobs 200 and 305 in May.
c. Prepare the journal entry on May 31 to apply factory overhead to both jobs in May according to the predetermined overhead rate. Refer to the Chart of Accounts for exact wording of account titles.

Chart of Accounts

CHART OF ACCOUNTS
Bergan Company
General Ledger
ASSETS
110 Cash
121 Accounts Receivable
125 Notes Receivable
126 Interest Receivable
131 Materials
132 Work in Process
133 Factory Overhead
134 Finished Goods
141 Supplies
142 Prepaid Insurance
143 Prepaid Expenses
181 Land
191 Factory
192 Accumulated Depreciation-Factory
LIABILITIES
210 Accounts Payable
221 Utilities Payable
231 Notes Payable
236 Interest Payable
241 Lease Payable
251 Wages Payable
252 Consultant Fees Payable
EQUITY
311 Common Stock
340 Retained Earnings
351 Dividends
390 Income Summary
REVENUE
410 Sales
610 Interest Revenue
EXPENSES
510 Cost of Goods Sold
520 Wages Expense
531 Selling Expenses
532 Insurance Expense
533 Utilities Expense
534 Office Supplies Expense
540 Administrative Expenses
560 Depreciation Expense-Factory
590 Miscellaneous Expense
710 Interest Expense

Factory Overhead

a. For Bergan Company, determine the predetermined factory overhead rate using direct labor hours as the activity base.

per direct labor hour

b. During May, Bergan Company accumulated 2,500 hours of direct labor costs on Job 200 and 3,000 hours on Job 305. Determine the amount of factory overhead applied to Jobs 200 and 305 in May.

Journal

c. Prepare the journal entry on May 31 to apply factory overhead to both jobs in May according to the predetermined overhead rate. Refer to the Chart of Accounts for exact wording of account titles.

PAGE 10

JOURNAL

DATE DESCRIPTION POST. REF. DEBIT CREDIT

1

2
Answer #1
a) Predetermined factory overhead rate = 620000/80000 = $            7.75 per DLH
b) Overhead applied to Job 200 = 2500*7.75 = $       19,375
Overhead applied to Job 305 = 3000*7.75 = $       23,250
Total overhead applied $       42,625
c) Work in process $       42,625
Factory overhead $     42,625